HUBBLE SPACE TELESCOPE DAILY REPORT #2604 PERIOD COVERED: 0000Z (UTC) 04/17/00 - 0000Z (UTC) 04/18/00 Daily Status Report as of 109/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 8460 (Cycle 8 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations completed with no reported problems. 1.2 Completed STIS/CCD 8439 (Bias Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no reported problems. 1.3 Completed STIS/CCD 8437 (Dark Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. The proposal completed nominally. 1.4 Completed Seven Sets of STIS/CCD 8414 (STIS Sparse Field CTE test {Cycle 8)) The Space Telescope Imaging Spectrograph (CCD) was used to conduct CTE measurements using the "sparse field test", along both the serial and parallel axes. This program needs special commanding to provide {a} off-center MSM positionings of some slits, and {b} the ability to read out with any amplifier {A, B, C, or D}. All exposures are internals. The observations were completed as planned, and no anomalies were reported. 1.5 Completed WF/PC-2 8248 (Observations of Potential JVAS/CLASS Gravitational Lenses) The WF/PC-2 was used to perform JVAS/CLASS gravitational lens surveys. These surveys have been the most successful so far at finding new arcsecond-scale gravitational lenses. A total of 17 confirmed systems have been discovered. One surprising result is the number of doubly-imaged lensed systems compared to "quad'' configurations. The doubles are under-represented in the current JVAS/CLASS lens sample by a factor 5 based on conventional models of the lensing galaxies. The observations completed with no reported anomalies. 1.6 Completed Six Sets of WF/PC-2 8442 (Cycle 8 Standard Darks) The WF/PC-2 was used to obtain dark calibration dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. Over an extended period these data will also provide a monitor of radiation damage to the CCDs. The proposal completed with no reported anomalies. 1.7 Completed STIS/CCD/MA1/MA2 8423 (FASTEX --- Faint Standard Extension) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to investigate fainter standard stars which are needed for the flux calibration of COS and for the prism modes on ACS. The proposal completed with no noted anomalies. 1.8 Completed Four Sets of WF/PC-2 8805 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the WFPC2 Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working. The late re-acquisition described in 2.1 occurred during the fourth iteration of this proposal, and may have affected some observations. Otherwise, the proposal completed with no further reported problems. 1.9 Completed WF/PC-2 8448 (Cycle 8 INTFLAT Sweeps, VISFLAT Mini-Sweep, and Linearity Test) The WFPC2 was used to monitor the pixel to pixel flatfield response and the VISFLAT lamp degradation, as well as detect any possible changes due to contamination. The observations completed nominally. 1.10 Completed STIS/CCD 8189 (The Source of Gamma Ray Bursts and the Nature of their Hosts) The Space Telescope Imaging Spectrograph (CCD) was used to make high resolution optical and ultraviolet observations of the Gamma Ray Burster (GRB) 991216 and its host galaxy. This comprehensive set of observations will provide new insights into the astrophysics of GRBs and their high redshift host galaxies. The observations were completed as planned, and no problems were reported. 1.11 Completed S/C 5582 (FOC/48 Monthly Activation (Camera Section Only): Cycle 4) This proposal was executed as scheduled, and no anomalies were noted. 1.12 Completed FOC/96 7923 (FOC F/96 Turn-on) The Faint Object Camera (f/96) was used to take an internal flat and a dark to ensure that the 30-day limit for the FOC is not exceeded. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed WF/PC-2 8459 (Decontaminations and Associated Observations Pt. 2/2) The WF/PC-2 was used to perform the monthly decons. Also included are instrument monitors tied to decons: photometric stability check, focus monitor, pre- and post-decon internals {bias, intflats, kspots, & darks}, and UV throughput checks. The observations were completed nominally. 1.14 Completed Two Sets of STIS/CCD 8549 (POMS Test Proposal: Non-Scripted Parallel Proposal Continuation I) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude, non-scripted parallel observations as part of a POMS test proposal. The observations were completed as planned, and no anomalies were reported. 1.15 Completed WF/PC-2 8301 (A Snapshot Survey of X-ray Selected Central Cluster Galaxies) The WF/PC-2 was used to make snapshot survey observations of the x-ray, central cluster galaxy A2294. Central cluster galaxies are the most massive stellar systems known and have been used as standard candles for many decades. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 6 Successful: 6 Scheduled Re-acquisitions: 9 Successful: 9 The re-acquisition at 108/161656Z required three attempts in order to achieve fine lock. The proposal outlined in 1.8 may have been affected. 2.2 FHST Updates: Scheduled: 15 Successful: 15 2.3 Operations Notes: The operations team remains in the Building 23 back-up facility. Using ROP SR-1A, the SSR EDAC error counter was cleared once. As directed by ROP DF-18A, the ESB was dumped and cleared at 108/2152Z. The NICMOS FPA temperature monitor for detector 3 was enabled at 108/2155Z per an operations request. The SSA transmitter was turned on and off at 109/0241Z per ROP IC-2. The ephemerides were uplinked at 109/0224Z, using ROP DF-07A. 3.0 SIGNIFICANT FORTHCOMING EVENTS: HST486 flight software 1.2E (TRSWCC/VTFE) is scheduled for uplink between 109/1035Z and 110/0100Z. Continuation of normal science observations and calibrations. /CAW